# The Additive Bachelier model with an application to the oil option market in the Covid period

### v2

@@ 3 Calibration scheme
We have already discussed about the separability property (11) of the Implied vol, that enables to separate the dependence due to the term-structure (controlled by the parameter $\sigma_{t}$) and the one deriving from the moneyness. A similar property can be rewritten in terms of $t$ and $\chi$.
@@
$d(\eta,k):=\sum_{i}\left[C^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta,k)\right]^{2}\;,$
selecting the optimal values of $\eta$ and kk. [^3] The model is calibrated on both calls and puts, considering only the OTM options, thanks to the call-put parity.
@@
A property that describes well the Implied vol surface in some markets (e.g. the equity and interest rate markets) is sticky delta. Sticky delta describes the situation where the Implied volatility remains unchanged (i.e., it sticks) with the moneyness when the underlying moves.
@@
Proof.
See Appendix B ∎
This model property is a great advantage from the practitioners’ perspective: hedges derived from this model remain stable if the market moves according to the sticky delta rule, i.e., both the volatility surface and prices (normalized with the discount) as functions of moneyness remain constant over time. This is the main difference wrt exponential models, where only the prices normalized by the ATM forward remain constant.
In the next Section, we’ll observe that while the model fits exactly the most liquid products (discount-rate, forward, and ATM vol), the other parameters ($\eta$ and $k$) remain relatively constant even in the high-volatile oil market of the first Covid months. Furthermore, to show this stability of prices with $\eta$ and $k$, we’ll also keep the same value of $\eta$ and $k$ even for the next day, obtaining prices that are indistinguishable from the calibrated prices with $\eta$ and $k$ optimally selected. We’ll repeat the valuation even keeping the same parameters ($\eta$ and $k$) for an entire week: observed price differences remain generally within 11-22 cents, well within bid-ask prices. [^4] A trader that keeps hedged his/her portfolio, in practice hedges exposures generated (according to a given model) by model parameters that –ideally– correspond to different risks (e.g., underlying, discount-rate, ATM vol, skew, etc…). If these parameters remain relatively constant over time when market conditions do not change significantly, he/she avoids paying transaction costs on trades due to exposures generated by an inadequate model description.
@@ 4 Calibration results
In Section 4.1, we describe the data used in the analysis; in Sections 4.2 and 4.3, we show the results of the calibration. In particular, in Section 4.2, we show the results for all the value dates considered in this analysis. In Section 4.3, we focus on a specific value date in the dataset, the $29^{\mathrm{th}}$ of April 2020 (one week after the announce from the CME 2020, that changed reference model into Bachelier), comparing our results with market data and the three benchmarks.
@@ 4.1 Dataset
The dataset is composed of daily market prices of call and put options on the WTI oil futures traded on the CME. We consider the whole period when the CME has switched the options pricing and valuation model to Bachelier (CME 2020). The time window of interest goes from the $23^{\mathrm{rd}}$ of April 2020 to the $31^{\mathrm{st}}$ of August 2020. Data are present in all weekdays with the exception of two public holidays in the USA: the $25^{\mathrm{th}}$ of May 2020 (Memorial Day) and the $3^{\mathrm{rd}}$ of July 2020 (Independence Day Holiday), when no options were traded. Daily prices are obtained from Bloomberg. The options are quoted for a grid of strikes that goes from 2.5 \$ to 245 \$. The grid size is equal to 0.5 \$ from 5 \$ to 146.5 \$, it’s equal to 2.5 \$ from 147.5 \$ to 195 \$, and it’s 5 \$ from 195 \$ to 245 \$. We consider all available liquid options’ expiry: from June 2020 (JUN20) to December 2022 (DEC22). We have considered as “risk-free” rate the Effective Federal Funds Rate OIS curve in the time window of interest up to the three years expiry. In Table 1, we report a summary statistics.
@@
From this dataset, we first replicate discount-rates and ATM forwards from option prices, as described in Azzone and Baviera (2021). To do so, we need the couples of call and put options traded for the same strike and expiry, for each value date. In Table 2, we show the number of couples of options available for each value date in the period of the analysis, we report mean, median, standard deviation (std), and quantiles (q) 5%, 95%.
@@
Table 2: We report the number of call-put options’ couples traded for each maturity and value date: mean, median, standard deviation (std), and quantiles (q) 5%, 95% considering each value date.
@@
In Table 3, we provide the descriptive statistics of the number of available OTM options, for all the strikes and value dates, we report mean, median, standard deviation (std), and quantiles (q) 5%, 95%.
@@
Table 3: Number of OTM options considered in the Volatility surface calibration. We report the number of OTM call ($K>F_{0}$) and put ($K<F_{0}$) in the moneyness range $30\,\mathdollar,30\,\mathdollar$]: mean, median, standard deviation (std), and quantiles (q) 5%, 95% considering each value date.
We observe from these Tables that the market in the period of interest (from April 2020 to August 2020) was relatively illiquid. We’ll notice in the following that, despite these market conditions, our model is able to describe adequately the key features of this market, with stable results across the value dates in the period of interest.
@@ 4.2 A glimpse on all value dates
In this Section, we present the results for all value dates in the dataset; in particular, we show the spread of options’ discount-rates on the OIS rates, and the parameters $\eta$ and $k$ of the Additive model; we present their evolution across the different value dates.
Figure 1: Spread of the discount-rates (from option prices) on the OIS curve, for all value dates considered in the analysis. The vertical lines correspond to the expiries of the first two options. In blue (dots and continuous line, left axis), we can see the spread corresponding to the first maturity: we observe that it rapidly increases when approaching options’ first expiry, reaching some hundreds of basis points. In green (squares and dashed line, right axis), we see the average of the spread corresponding to the other expiries: we can observe that it is of few basis points (an average of 22 bps), i.e. its order of magnitude is much smaller than the spread for the first expiry.
Figure 1 shows the spread of the discount-rates derived from option prices on the “risk-free” rates (OIS) for different value dates. In the blue line (dots and continuous line, left axis), we can see the spread corresponding to the first maturity: this quantity shows a rapid increase when approaching the first expiry. The average of the corresponding spread on all the other expiries is plotted in green (squares and dashed line, right axis). We observe that the latter remains of the order of few basis points (on average around $2$ bps), i.e. negligible for all practical purposes in the volatile market of the first Covid period, and it is much smaller than the spread on the first expiry (even of $2$ orders of magnitude), that reaches some hundreds of basis points. The inability to reproduce even OIS rates is a clear indicator of market distress. The observed behaviour of option discount-rates for very short-term maturities highlights the extremely tough market conditions that the WTI oil market experienced during the Covid period.
For this reason, the options on the first expiry are not considered in the calibration procedure when the discount-rates are significantly different from the “risk-free” rates; in particular, the adopted criterion is to consider the first expiry when the spread is up to one order of magnitude larger than the average spread (i.e., it is less than $20$ bps).
@@
Figure 3 shows the behaviour of the parameter $k$ for the different value dates during the analysed period. The level of $k$ remains close to one up to mid July where it reached –moving slowly over time– a level close to $0.5$ where it remained up to the end of August indicating relative calmer period of the Covid-19 pandemic during the summer time.
Figure 2: Parameter $\eta$ of the Additive Bachelier calibrated for the different value dates considered in the analysis. The vertical lines correspond to the expiries of the first two options. We observe that its value does change significantly for the value dates, but it moves slowly over time.
Figure 3: Parameter $k$ of the Additive Bachelier calibrated for the different value dates considered in the analysis. The vertical lines correspond to the expiries of the first two options. We observe that its value is rather close to one and moves slowly over time.
@@ 4.2.1 Quality of calibration
$RMSE_{t_{0}}(\eta^{*},k^{*}):=\sqrt{\frac{1}{N_{t_{0}}}\sum_{i}\left[C_{t_{0}}^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta^{*},k^{*})\right]^{2}}\;,$
@@
As shown in Figure 4, the increase in error when using the previous value date’s parameters is contained and always in the order of one cent or below. This fact suggests that the model parameters $\eta,\,k$ are stable over two consecutive value dates, i.e., their small deviations do not significantly worsen the model accuracy.
Figure 4: For each value date $t_{0}$, root mean square error between market prices and model prices using the parameters $\eta_{t_{0}},\,k_{t_{0}}$ calibrated for the value date $t_{0}$, $RMSE_{t_{0}}(\eta_{t_{0}},k_{t_{0}})$ , and the RMSE between market prices and model prices obtained using the previous day parameters $\eta_{t_{0}-1},\,k_{t_{0}-1}$ $RMSE_{t_{0}}(\eta_{t_{0}-1},k_{t_{0}-1})$. The difference between the two errors is negligible for almost all dates (of the order of 11 cent or below), confirming the stability over two consecutive value dates of the Additive Bachelier model.
@@
$RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-4:t_{0}},k_{t_{0}-4:t_{0}}):=\sqrt{\frac{1}{\sum_{s=t_{0}-4}^{t_{0}}N_{s}}\sum_{s=t_{0}-4}^{t_{0}}\sum_{i}\left[C_{s}^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta_{s},k_{s})\right]^{2}}\;,$
@@
$RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-5},k_{t_{0}-5}):=\sqrt{\frac{1}{\sum_{s=t_{0}-4}^{t_{0}}N_{s}}\sum_{s=t_{0}-4}^{t_{0}}\sum_{i}\left[C_{s}^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta_{t_{0}-5},k_{t_{0}-5})\right]^{2}}\;,$
@@
As already mentioned commenting Tables 2 and 3, the WTI derivative market was relatively illiquid in the time window of interest. We highlight that, despite the observed high volatility and low liquidity market, it is possible to achieve stable results over time in the calibration of the Additive Bachelier parameters ($\eta$ and $k$). In the following Section, we discuss the calibration results achieved on a single value date.
@@
Figure 5: Difference of RMSE over a 55-day window using fixed versus rolling parameters. We compare the 55-day RMSE between market prices and model prices using the parameters $\eta_{s},\,k_{s}$ for the each value date $s\in\left[t_{0}-4,\,t_{0}\right]$, $RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-4:t_{0}},k_{t_{0}-4:t_{0}})$, against the $RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-5},k_{t_{0}-5})$ obtained using the fixed parameters $\eta_{t_{0}-5},\,k_{t_{0}-5}$, calibrated before the beginning of the window. We observe a contained increase in the calibration error of about 11 cent or below, with the only exception of the $27^{\mathrm{th}}$ of May 2020, when it is less than $2.5$ cent. This fact highlights the robustness of the Additive Bachelier model’s parameters.
@@ 4.3 Detailed results for one date: the $\mathbf{29}^{\mathbf{th}}$ of April 2020
In this Section, we show the results with value date on the $29^{\mathrm{th}}$ of April 2020 (one week after the announce from the CME 2020, in one of the most volatile periods), [^5] Results do not appear significantly different on all other value dates and they are available upon request. compared with the results of other three models considered as benchmarks:
  • the Lévy Bachelier model (hereinafter “Lévy model”), defined in (21), Appendix A,
  • a Lévy Bachelier model that includes options with one ttm at a time (named “slice-by-slice”),
  • an Additive Logistic model (Carr and Torricelli 2021).
@@
As already mentioned, since the spread on the first date is larger than $20$ basis points, we do not consider the first expiry in the next steps of the calibration.
@@
Figure 7: ATM volatility comparison between market data (red, squares and continuous line), the Additive model (in blue, dots and continuous line), the slice-by-slice (green, dots dashed line), the Lévy model (black, dashed line) and the Additive Logistic (cyan, continuous line), on the $29^{\mathrm{th}}$ of April 2020, for $\alpha=1/2$. While the first three are indistinguishable in the Figure with both the slice-by-slice and the Additive Bachelier replicating exactly market ATM vol, both the Lévy model and the Additive Logistic are not able to reproduce the ATM vol.
@@
We compare the results for the four models observing the root mean square error (RMSE) of model prices wrt market prices separately for each maturity $t$ (with moneyness in the interval $30\,\mathdollar,30\,\mathdollar$]). We plot in Figure 9 the RMSE from each model for all the expiries (except for the first one) present in the dataset.
@@
We observe that the error from the Lévy model and the Additive Logistic are almost one order of magnitude (two for longer expiries in the case of the Lévy model) larger than the RMSE of the slice-by-slice, while the error of the Additive Bachelier (in blue, dots and continuous line) is quite close to the one of the slice-by-slice for expiries up to one year, becoming larger for longer expiries.

### v3

@@ 3 Calibration scheme
We have already discussed the separability property (12) of the implied vol, which enables us to separate the dependence due to the term-structure (controlled by the parameter $\sigma_{t}$) and the one deriving from the moneyness. A similar property can be rewritten in terms of $t$ and $\chi$.
@@
In the marketplace, it is now standard to obtain ATM forwards and the discount factors via a simple linear regression of call/put prices, following the methodology described in Azzone and Baviera (2021). Specifically, for each maturity, a linear regression of the difference between observed call and put prices against their strike prices is performed. This allows us to identify the market-implied discount factor and the ATM forward price simultaneously. This step ensures that the subsequent calibration of the volatility surface is consistent (and coincides exactly) with the market-implied term structures of forwards.
@@
$d(\eta,k):$ $=\sum_{i}\left[C^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta,k)\right]^{2}=$ (20)
${\color[rgb]{0,0,0}=\sum_{i}\left[C^{mkt}(\chi_{i}\,\sigma_{t_{i}}^{ATM}\,\sqrt{t_{i}},t_{i})-B_{0}\,\sigma_{t_{i}}^{ATM}\,\sqrt{t_{i}}\,\pazocal{C}({\chi}_{i};\eta,k)\right]^{2}}\;,$
selecting the optimal values of $\eta$ and kk. [^8] The model is calibrated on both calls and puts, considering only the OTM options, thanks to the call-put parity. Let us emphasize, that the evaluation of the distance $d(\eta,k)$ in (20) is particularly fast for this model: it requires the computation of only $\pazocal{C}(\chi,\eta,k)$ for different values of $\chi_{i}$ that can be obtained applying only one fast Fourier transform.
@@
A property that describes well the implied vol surface in some markets is sticky delta. Sticky delta describes the situation where the implied volatility remains unchanged (i.e. it sticks) with the moneyness when the underlying moves.
@@
This model property is a great advantage from the practitioners’ perspective: hedges derived from this model remain stable if the market moves according to the sticky delta rule, i.e. both the volatility surface and prices (normalized with the discount) as functions of moneyness remain constant over time. This is another difference wrt exponential models, where only the prices normalized by the ATM forward remain constant.
In the next Section, we’ll observe that while the model fits exactly the most liquid products (discount-rate, forward, and ATM vol), the other parameters ($\eta$ and $k$) remain relatively constant even in the high-volatile oil market of the first Covid months. Furthermore, to show this stability of prices with $\eta$ and $k$, we’ll also keep the same value of $\eta$ and $k$ even for the next day, obtaining prices that are almost indistinguishable from the calibrated prices with $\eta$ and $k$ optimally selected. We’ll also repeat the valuation even keeping the same parameters ($\eta$ and $k$) for an entire week: observed price differences remain generally within 11-22 cents, well within bid-ask prices. [^9] A trader that keeps hedged his/her portfolio, in practice hedges exposures generated (according to a given model) by model parameters that –ideally– correspond to different risks (e.g., underlying, discount-rate, ATM vol, skew, etc…). If these parameters remain relatively constant over time when market conditions do not change significantly, he/she avoids paying transaction costs on trades due to exposures generated by an inadequate model description.
@@ 4 Calibration results
In Section 4.1, we describe the data used in the analysis; in Sections 4.2 and 4.3, we show the results of the calibration. In particular, in Section 4.2, we show the results for all the value dates considered in this analysis. In Section 4.3, we focus on a specific value date in the dataset, the $29^{\mathrm{th}}$ of April 2020 (one week after the announcement from the CME, that changed the reference model into Bachelier), comparing our results with market data and three benchmarks.
@@ 4.1 Dataset
The dataset is composed of daily market prices of call and put options on the WTI oil futures traded on the CME. We consider the whole period when the CME has switched the options pricing and valuation model to Bachelier (CME 2020). The time window of interest goes from the $23^{\mathrm{rd}}$ of April 2020 to the $31^{\mathrm{st}}$ of August 2020. Data are present on all weekdays except two public holidays in the USA: the $25^{\mathrm{th}}$ of May 2020 (Memorial Day) and the $3^{\mathrm{rd}}$ of July 2020 (Independence Day holiday), when no options were traded. Daily prices are obtained from Bloomberg. The options are quoted for a grid of strikes that goes from 2.5 \$ to 245 \$. The grid size is equal to 0.5 \$ from 5 \$ to 146.5 \$, it’s equal to 2.5 \$ from 147.5 \$ to 195 \$, and it’s 5 \$ from 195 \$ to 245 \$. We consider all available liquid options’ expiry: from June 2020 (JUN20) to December 2022 (DEC22). We have considered as “risk-free” rate the Effective Federal Funds Rate OIS curve in the time window of interest up to the three years expiry. In Table 1, we report a summary statistics.
@@
From this dataset, we first replicate discount-rates and ATM forwards from option prices, as described in Azzone and Baviera (2021). To do so, we need the couples of call and put options traded for the same strike and expiry, for each value date. In Table 2, we show the number of couples of options available for each value date in the period of the analysis. We report mean, median, standard deviation (std), and quantiles (q) 5%, 95%.
@@
Table 2: We report the number of call-put options’ couples traded for each maturity and value date: mean, median, standard deviation (std), and quantiles (q) 5%, 95%, considering each value date.
@@
In Table 3, we provide the descriptive statistics of the number of available OTM options, for all the strikes and value dates. We report mean, median, standard deviation (std), and quantiles (q) 5%, 95%.
@@
Table 3: Number of OTM options considered in the volatility surface calibration. We report the number of OTM calls ($K>F_{0}$) and puts ($K<F_{0}$) in the moneyness range $30\,\mathdollar,30\,\mathdollar$]: mean, median, standard deviation (std), and quantiles (q) 5%, 95% considering each value date.
We observe from these Tables that the market in the period of interest (from April 2020 to August 2020) was relatively illiquid. We’ll notice in the following that, despite these market conditions, our model is able to describe the key features of this market adequately, with stable results across the value dates in the period of interest.
@@ 4.2 A glimpse on all value dates
In this Section, we present the results for all value dates in the dataset; in particular, we show the spread of options’ discount-rates on the OIS rates, and the parameters $\eta$ and $k$ of the additive model; we present their evolution across the different value dates. In this Section, we consider the model with the choice of $\alpha=1/2$ (results for other values of $\alpha$ are similar).
Figure 1: Spread of the discount-rates (from option prices) on the OIS curve, for all value dates considered in the analysis. The vertical lines correspond to the expiries of the first two options. In blue (dots and continuous line, left axis), we can see the spread corresponding to the first maturity: we observe that it rapidly increases when approaching options’ first expiry, reaching some hundreds of basis points. In green (squares and dashed line, right axis), we see the average of the spread corresponding to the other expiries: we can observe that it is of a few basis points (an average of 22 bps), i.e. its order of magnitude is much smaller than the spread for the first expiry.
Figure 1 shows the spread of the discount-rates derived from option prices on the “risk-free” rates (OIS) for different value dates. In the blue line (dots and continuous line, left axis), we can see the spread corresponding to the first maturity: this quantity shows a rapid increase when approaching the first expiry. The average of the corresponding spread on all the other expiries is plotted in green (squares and dashed line, right axis). We observe that the latter remains of the order of a few basis points (on average around $2$ bps), i.e. negligible for all practical purposes in the volatile market of the first Covid period, and it is much smaller than the spread on the first expiry (even of $2$ orders of magnitude), that reaches some hundreds of basis points. The inability to reproduce even OIS rates is a clear indicator of market distress. The observed behaviour of option discount-rates for very short-term maturities highlights the extremely tough market conditions that the WTI oil market experienced during the Covid period.
For this reason, the options on the first expiry are not considered in the calibration procedure when the discount-rates are significantly different from the “risk-free” rates; in particular, the adopted criterion is to consider the first expiry when the spread is up to one order of magnitude larger than the average spread (i.e. it is less than $20$ bps).
@@
Figure 3 shows the behaviour of the parameter $k$ for the different value dates during the analysed period. The level of $k$ remains close to one up to mid July, where it reached –moving slowly over time– a level close to $0.5$ where it remained up to the end of August, indicating relative calmer period of the Covid-19 pandemic during the summer time.
Figure 2: Parameter $\eta$ of the additive Bachelier calibrated for the different value dates considered in the analysis. The vertical lines correspond to the expiries of the first two options. We observe that its value does not change significantly for the value dates, and it moves slowly over time.
Figure 3: Parameter $k$ of the additive Bachelier calibrated for the different value dates considered in the analysis. The vertical lines correspond to the expiries of the first two options. We observe that its value moves slowly over time.
@@ 4.2.1 Quality of calibration
$RMSE_{t_{0}}(\eta^{*},k^{*}):=\sqrt{\frac{1}{N_{t_{0}}}\sum_{i=1}^{N_{t_{0}}}\left[C_{t_{0}}^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta^{*},k^{*})\right]^{2}}\;,$
@@
As shown in Figure 4, the increase in error when using the previous value date’s parameters is contained and always in the order of one cent or below. This fact suggests that the model parameters $\eta,\,k$ are stable over two consecutive value dates, i.e. their small deviations do not significantly worsen the model accuracy.
Figure 4: For each value date $t_{0}$, RMSE between market prices and model prices using the parameters $\eta_{t_{0}},\,k_{t_{0}}$ calibrated for the value date $t_{0}$, $RMSE_{t_{0}}(\eta_{t_{0}},k_{t_{0}})$, and the RMSE between market prices and model prices obtained using the previous day parameters $\eta_{t_{0}-1},\,k_{t_{0}-1}$ $RMSE_{t_{0}}(\eta_{t_{0}-1},k_{t_{0}-1})$. The difference between the two errors is negligible for almost all dates (of the order of 11 cent or below), confirming the stability over two consecutive value dates of the additive Bachelier model.
@@
$RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-4:t_{0}},k_{t_{0}-4:t_{0}}):=\sqrt{\frac{1}{\sum_{s=t_{0}-4}^{t_{0}}N_{s}}\sum_{s=t_{0}-4}^{t_{0}}\sum_{i=1}^{N_{s}}\left[C_{s}^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta_{s},k_{s})\right]^{2}}\;,$
@@
$RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-5},k_{t_{0}-5}):=\sqrt{\frac{1}{\sum_{s=t_{0}-4}^{t_{0}}N_{s}}\sum_{s=t_{0}-4}^{t_{0}}\sum_{i=1}^{N_{s}}\left[C_{s}^{mkt}(x_{i},t_{i})-C^{mod}(x_{i},t_{i};\eta_{t_{0}-5},k_{t_{0}-5})\right]^{2}}\;,$
@@
As already mentioned, commenting on Tables 2 and 3, the WTI derivative market was relatively illiquid in the time window of interest. We highlight that, despite the observed high volatility and low liquidity market, it is possible to achieve stable results over time in the calibration of the additive Bachelier parameters ($\eta$ and $k$). In the following Section, we discuss the calibration results achieved on a single value date.
@@
Figure 5: Difference of RMSE over a 55-day window using fixed versus optimal parameters. We compare the 55-day RMSE between market prices and model prices using the parameters $\eta_{s},\,k_{s}$ for the each value date $s\in\left[t_{0}-4,\,t_{0}\right]$, $RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-4:t_{0}},k_{t_{0}-4:t_{0}})$, against the $RMSE_{t_{0}-4:t_{0}}(\eta_{t_{0}-5},k_{t_{0}-5})$ obtained using the fixed parameters $\eta_{t_{0}-5},\,k_{t_{0}-5}$, calibrated before the beginning of the window. We observe a contained increase in the calibration error of about 11 cent or below, with the only exception of the $27^{\mathrm{th}}$ of May 2020, when it is less than $2.5$ cents. This fact highlights the robustness of the additive Bachelier model’s parameters.
@@ 4.3 Detailed results for one date: the $\mathbf{29}^{\mathbf{th}}$ of April 2020
In this Section, we show the results with value date on the $29^{\mathrm{th}}$ of April 2020 (one week after the announcement from the CME, in one of the most volatile periods), [^10] Results do not appear significantly different on all other value dates and they are available upon request. compared with the results of other three models considered as benchmarks:
@@
As already mentioned, since the spread on the first date is larger than $20$ bps, we do not consider the first expiry in the next steps of the calibration.
@@
Figure 7: ATM volatility comparison between market data (red, squares and continuous line), the additive model (in blue, dots and continuous line), the slice-by-slice (green, dots dashed line), the Lévy model (black, dashed line) and the additive logistic (cyan, continuous line), on the $29^{\mathrm{th}}$ of April 2020, for $\alpha=1/2$. While the first three are indistinguishable in the Figure, with both the slice-by-slice and the additive Bachelier replicating exactly the market ATM vol, both the Lévy model and the additive logistic are not able to reproduce the ATM vol.
@@
We compare the results for the four models observing the RMSE of model prices wrt market prices separately for each maturity $t$ (with moneyness in the interval $30\,\mathdollar,30\,\mathdollar$]). We plot in Figure 9 the RMSE from each model for all the expiries (except for the first one) present in the dataset.
@@
We observe that the error from the Lévy model and the additive logistic are almost one order of magnitude larger than the RMSE of the slice-by-slice, while the error of the additive Bachelier (in blue, dots and continuous line) is quite close to the one of the slice-by-slice for expiries up to one year, becoming larger for longer expiries.